OpenStudy (anonymous):

Which function is a linear function? 1 − 3x2 = -y y + 7 = 5x x3 + 4 = y 9(x2 − y) = 3 y − x3 = 8

OpenStudy (danjs):

highest power of 1 on the independent variableeee

OpenStudy (anonymous):

huh

OpenStudy (danjs):

only one has a power of 1 , the others are polynomials, not linear

OpenStudy (anonymous):

i know i think

OpenStudy (anonymous):

so its C

OpenStudy (cmtboy2):

x3 + 4 = y

OpenStudy (anonymous):

yeah it's c

OpenStudy (danjs):

y = x^3 + 4 is a polynomial, not linear

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

should a linear function start with Y

OpenStudy (anonymous):

shouldn't it?

OpenStudy (danjs):

remember a line has a form Ax+By=C , or y = m*x+b, the highest power on x or the y is 1, no larger

OpenStudy (anonymous):

wat are examples of nonlinear

OpenStudy (danjs):

polynomials are y = ax^n + bx^(n-1) ..... c^x + d powers larger than 1 are not lines

OpenStudy (anonymous):

ight here is a nother ?

OpenStudy (danjs):

notice the second choice is the only one that has a highest power of 1, the others are greater

OpenStudy (anonymous):

In the coming year, a vehicle manufacturer has decided to manufacture 150 vehicles per day. The function v = 150d represents the company’s production for the coming year, v, with respect to the number of days, d. The rate of change of the function representing the number of vehicles manufactured for the coming year is , and its graph is a . So, the function is a function.

OpenStudy (danjs):

if a function has a constant rate of change (slope) all the time, it is a line

OpenStudy (danjs):

notice as before, the highest power is 1 v = 150*d linear y = m*x + b, rate of change is the slope,m

OpenStudy (anonymous):

is it a straight line

OpenStudy (anonymous):

@DanJS